OpenStudy (anonymous):

Add or subtract as indicated. 5/X-2-3/X-2

OpenStudy (anonymous):

is it \[\frac{ 5 }{ x }-2-\frac{ 3 }{ x }-2\]

OpenStudy (anonymous):

@SADIQ6205

OpenStudy (anonymous):

thx for helping me um gimme a sex

OpenStudy (anonymous):

no its 5/(x-2)-3/(x-2)

OpenStudy (anonymous):

What do u mean by sex?

OpenStudy (anonymous):

sorry i meant sec...

OpenStudy (anonymous):

be careful next time

OpenStudy (anonymous):

lol sorry thx :) did u take offense, i didnt mean it.

OpenStudy (anonymous):

would u have minded if i said it, it could have been an accident

OpenStudy (anonymous):

um hullo

OpenStudy (anonymous):

@sheelo_mughal hey are u still helpiong me?

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

yes actually my laptop was corrupting so i restarted it

OpenStudy (anonymous):

oh okay i though u got mad and left so whenever ur ready u can start helping

OpenStudy (anonymous):

can u tell me exact answer?

OpenStudy (anonymous):

actually i got that question can u help me with another

OpenStudy (anonymous):

i mean u must have answers in ur book

OpenStudy (anonymous):

no i'll help with this first

OpenStudy (anonymous):

but i solved it already it was simple, i just overlooked an easy part of the question

OpenStudy (anonymous):

ok so what next one :/

OpenStudy (anonymous):

here it is thx so much

OpenStudy (anonymous):

thx guys, i thinks its easy idk what to do

OpenStudy (anonymous):

so im weak in graphs

OpenStudy (anonymous):

i mean sorry

OpenStudy (anonymous):

oh okay no problem, i hope brainy beauty can help

OpenStudy (anonymous):

\[= \frac{ 5-3 }{ x-2}\]

OpenStudy (anonymous):

not the tope problem, i got that thx, but the one that comes down to the bottom of the convo the graph one? sorry for the confusion

OpenStudy (anonymous):

= 2/ (x-2)

OpenStudy (anonymous):

Okay wait

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

sorry but dont go that personal

OpenStudy (anonymous):

hey who's this in ur profile pic

OpenStudy (anonymous):

u dont know him?

OpenStudy (anonymous):

looks like he's old indian actor

OpenStudy (anonymous):

yep lol, i love his acting, im 16 btw, not getting personal just fyi

OpenStudy (anonymous):

@sheelo_mughal

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so u dont know how to do it?

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

It should be a. Take any point on the graph. Put the value of x in every equation. If you get the answer as y, that option is correct. Try it.

OpenStudy (anonymous):

okay, um gimem a sec let me try it

OpenStudy (anonymous):

Yeah sure

OpenStudy (anonymous):

@SADIQ6205 r u indian

OpenStudy (anonymous):

got it its a and yeaah obviously

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@sheelo_mughal and u must be african? :P

OpenStudy (anonymous):

thx so much brain beauty

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You're welcome! :)

OpenStudy (anonymous):

@sheelo_mughal i was kidding

OpenStudy (anonymous):

@sheelo_mughal this is why u need to learn to make friends and accept text requests :)

OpenStudy (anonymous):

ok first of all whats ur religion

OpenStudy (anonymous):

whats urs?

OpenStudy (anonymous):

im muslim...

OpenStudy (anonymous):

sikh and hindu...

OpenStudy (anonymous):

jk im muslim and im pretty sure ur hindu right?

OpenStudy (anonymous):

me too muslim And i thank Allah for this Alhamdulillah

OpenStudy (anonymous):

dont think of anybody whether he/she is sikh or hindu just pray to Allah for them to become muslim

OpenStudy (anonymous):

@sheelo_mughal sorry my internet crashed and ur so right i was actually jsut messing with u

OpenStudy (anonymous):

and alhamdulillah ur muslim, u sunni or shia?